### v1

@@ A.3 Basics from algebraic number theory
B.2 Proof of Theorem 1.4
We have proved in Section B.1 that if the Fourier weight of the function $f$ is concentrated on the $s$-largest Fourier coefficients, then $f$ is $\frac{\mu}{\sqrt{s}}$ close to a $\lceil\log_{p}s\rceil+1$-granular in the $\ell_{2}$ norm. In this subsection, we will prove Theorem 1.4, which says that this function must be a Boolean valued function.
Proof B.5.
Let us assume $k=\lceil\log_{p}s\rceil+1$. Let $B$ be the set of character functions corresponding to the $s$ largest coefficients and $S$ be set of the rest of the characters. Then, by Theorem 1.3, for all $\chi_{r_{i}}\in B$, $\widehat{f}(\chi_{r_{i}})$ is $\frac{\mu}{\sqrt{s}}-$close to $\lceil\log_{p}s\rceil+1$-granular. Therefore, for all $\chi_{r_{i}}\in B$, we can write $f(\chi_{r_{i}})$ as the sum of $\widehat{F}(\chi_{r_{i}})$ and $\widehat{G}(\chi_{r_{i}})$, that is,
@@
where $|\widehat{F}(\chi_{r_{i}})|$ is $\lceil\log_{p}s\rceil+1$-granular and $|\widehat{G}(\chi_{r_{i}})|\leq\frac{\mu}{\sqrt{s}}$.
@@
Clearly, $F$ is $s-$sparse and $|\widehat{F}(\chi)|$ is $\lceil\log_{p}s\rceil+1$-granular for all Fourier coefficients $\widehat{F}(\chi)$ of $F$.
@@
$1=f^{2}=(F+G)^{2}=F^{2}+2FG+G^{2}=F^{2}+G(2F+G)=F^{2}+G(2f-G).$ (B.6)
Let $H=G(2f-G)$. Let us look at the Fourier coefficients $\widehat{H}(\chi)$ of $H$.
$|\widehat{H}(\chi)|$ $\leq||G||_{2}||2f-G||_{2}$ by Lemma A.12
$\leq||G||_{2}(||2f||_{2}+||G||_{2})$
@@
$\leq\frac{1}{2(p^{2}s)^{p-1}}.$ (B.7)
Now, since $\widehat{F}(\chi)$ is $\lceil\log_{p}s\rceil+1$-granular, any Fourier coefficient $\widehat{F^{2}}(\chi)$ of $F^{2}$ is $22\lceil\log_{p}s\rceil+$ 2-granular. That is, $\widehat{F^{2}}(\chi)=\frac{g^{\prime}(\omega_{p})}{p^{2k}}$, where $k=\lceil\log_{p}s\rceil+1$ and $g^{\prime}(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq p-1$. Also, by Remark 2.3, $|\widehat{F^{2}}(\chi)|\geq\frac{1}{(p^{2}s)^{p-1}}$.
Now, from Equation B.6, we have,
@@
where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathbb{Z}_{p}^{n}$, as defined in Lemma A.2, part 1.
Since $\widehat{F^{2}}(\chi)\geq\frac{1}{(p^{2}s)^{p-1}}$ and $\widehat{H}(\chi)<\frac{1}{2(p^{2}s)^{p-1}}$, these two cannot sum up to 00; therefore, $\widehat{F^{2}}(\chi)=0$, $\widehat{H}(\chi)=0$ for all $\chi\neq 0$.
@@
[…]
$\Rightarrow$ $\,\widehat{H}(\chi_{0})\;=\;\frac{p^{2k}\,-\,g^{\prime}(\omega_{p})}{p^{2k}}\;,$
which implies that $\widehat{H}(\chi_{0})$ is $2k$-granular if $\widehat{H}(\chi_{0})\neq 0$. So, by Remark 2.3, $|\widehat{H}(\chi_{0})|\geq\frac{1}{(p^{2}s)^{p-1}}$, which is not possible since $|\widehat{H}(\chi_{0})|\leq\frac{1}{2(p^{2}s)^{p-1}}$ by B.6. So, $\widehat{H}(\chi_{0})=0$, hence $\widehat{F^{2}}(\chi_{0})=1$. Therefore $F^{2}(x)=0$ for all $x\in\mathbb{Z}_{p}^{n}$, which implies that the range of $F$ is $\{-1,+1\}$.
@@
Appendix C A $p$-independent lower bound
In this section, we discuss an interesting property of Boolean-valued functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, where $p$ is a prime number, and eventually establish a $p$-independent lower bound on the Fourier coefficient of an $s$-sparse Boolean-valued function.
For $x\in\mathbb{Z}_{p}^{n}$ and $i\in[p-1]$, we define $i^{-1}x:=(i^{-1}x_{1},\ldots,i^{-1}x_{n})$, where $x=(x_{1},\ldots,x_{n})$ and $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{p}^{*}$ of $\mathbb{Z}_{p}$.
Lemma C.1.
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, and $\chi\in\operatorname{supp}(f)$. Then, for each $i\in\{1,2,\ldots,p-1\}$, there exists a function $h_{i}:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, defined by $h_{i}(x)=f(i^{-1}x)$ such that
@@ Appendix C A $p$-independent lower bound
where $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{p}^{*}$ of $\mathbb{Z}_{p}$.
Proof C.2.
Proof of Part (1). Let $h_{i}(x):=f(i^{-1}x)$. Note that, by definition, $h_{i}$ is a Boolean-valued function. Further, by Definition A.5:
$\hat{h_{i}}(\chi)$ $=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}h_{i}(x)\overline{\chi(x)}$
$=\frac{1}{p^{n}}\sum_{ix\in\mathbb{Z}_{p}^{n}}h_{i}(ix)\overline{\chi(ix)}$
$=\frac{1}{p^{n}}\sum_{ix\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)^{i}}$
$=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}$
@@
Proof C.3.
By Definition A.5, we have $\hat{f}(\chi)=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}=\frac{u(\omega_{p})}{p^{n}}$, where $u(\omega_{p})=\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}$. Think of $u(X)\in\mathbb{Z}[X]$. Then, $\hat{h_{i}}(\chi)=\hat{f}(\chi^{i})=\frac{1}{p^{n}}\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}=\frac{u(\omega_{p}^{i})}{p^{n}}$; the last holds since
$u(\omega_{p})\;=\;\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}\;\iff\;u(\omega_{p}^{i})\;=\;\sum_{x\in\mathbb{Z}_{p}^{n}}f(x)\overline{\chi(x)}^{i}\;.$
Using Fact 1, we know that $u(\omega_{p})\neq 0\iff u(\omega_{p}^{i})\neq 0$. Hence, whenever $f(\chi)\neq 0,\ h_{i}(\chi)\neq 0$ as well, and the vice versa. Thus, the claim holds.
The following corollary follows immediately from Lemma C.1 (part 2).
Corollary C.4.
@@
The following corollary follows immediately from Lemma C.1.
Corollary C.6.
For all function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, $s_{f}\geq p-1$.
Proof C.7.
From the proof of Lemma C.1, it is clear that if $\chi\in\operatorname{supp}(f)$, then $\chi^{i}\in\operatorname{supp}(f)$ for all $i\in\{1,\ldots,p-1\}$. Hence, sparsity of $f,\ s_{f}\geq p-1$.
Now, we prove a lower bound on the Fourier coefficient which is $p$-independent; this is an immediate consequence of Corollary C.6.
Theorem C.8 ($p$-independent lower bound).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a function. Let $\chi\in\sup(f)$. Then,
@@
Proof C.9.
We know that $s_{f}\geq p-1$, by Corollary C.6. Hence, we combine this with Theorem 1.5 , to get
$|\widehat{f}(\chi)|\;\geq\;\frac{1}{(p^{2}s_{f})^{(p-1)/2}}\;>\;\frac{1}{((s_{f}+1)\sqrt{s_{f}})^{s_{f}}};.$
In the above, we used the fact that $(p^{2}s_{f})^{1/2}=p\sqrt{s_{f}}\leq(s_{f}+1)\sqrt{s_{f}}$.
Appendix D Proofs of Lemma 3.2 and Lemma 3.4
In this section, we prove Lemma 3.2 and Lemma 3.4, which are stated in Section 3.
D.1 Proof of Lemma 3.2
Proof D.1 (Proof of Lemma 3.2).
@@
where we simply used the GP series with the ratio $-\omega_{p}$. Note that
@@
[…]
@@
[…]
Note that, for large enough $n$, and a fixed $p\geq 5$, $c>1$. Hence, the proof follows.
D.2 Proof of Lemma 3.4
Proof D.2 (Proof of Lemma 3.4).
@@
So comparing with Equation D.1, we can see that $c$ is bigger in this case.
Remark D.3.
When $p=3$, we know that $12\cos(\pi/3)=$ 1, and hence $c<1$ ($c\to 1$, when $n\to\infty$).
Remark D.4.
@@
$c=\log_{5}|(2\cos(\frac{\pi}{5}))|+1-\frac{1}{2}\cdot\log_{5}(2)=1.265...$ (D.1)
@@
Appendix E Testing $s$-sparsity
In this section our goal is to design an algorithm that given a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ accepts (with probability $\geq\frac{2}{3}$) if $f$ is $s$-sparse and rejects (with probability $\geq\frac{2}{3}$) if $f$ is $\epsilon$-far from any $s$-sparse Boolean function, where $p$ is an odd prime.
The algorithm’s primary concept closely aligns with that presented in [15]. Since we know from Theorem 1.5 that all the coefficients of an $s$-sparse function are more than $1/(p^{2}s)^{\lceil(p-1)/2\rceil}$, so the idea is to partition the Dual group (the set of characters) into $\Theta(s^{2})$ buckets, and for each bucket $B$ estimate the weight $\mathsf{wt}(B)$ of the bucket which is defined as follows:
Definition E.1 (The weight function).
Let $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ be a function. The weight of a bucket $B$ is defined by

### v2

@@ 3.1 The projection on a coset
3.3 Lower bound on the Fourier coefficients
In this section, we will prove Theorem 1.5 assuming Theorem 1.3, which gives us a lower bound on the Fourier coefficients of functions from $\mathcal{G}$ to $\{-1,+1\}$. This is a generalization on the granularity of a function $f:\mathbb{Z}_{2}^{n}\to\mathbb{R}$ when the domain of the function is $\mathcal{G}$. We will use the following theorem; for more details see Section 2.4.
Theorem 3.8.
$\bigg{|}\prod_{i\in\mathbb{Z}_{n}^{\ast}}f(\omega_{n}^{i})\bigg{|}\;\geq\;1\;.$
Proof of Theorem 1.5.
If we put $\mu=0$ in Theorem 1.3, we get that there exists a $g\in\mathbb{Z}[X]$, such that $|\widehat{f}(\chi)|\geq|g(\omega_{\Delta})/\Delta^{k}|$, where $k=\lceil\frac{\log s}{4(\log\Delta-\log\varphi(\Delta))}-1\rceil$, $s_{f}$ being the sparsity of the Boolean valued function $f$. We know by Theorem 3.8 that $|\prod_{i\in\mathbb{Z}_{\Delta}^{\times}}g(\omega_{\Delta}^{i})|\geq 1$. Therefore,
$\bigg{|}\prod_{i\in\mathbb{Z}_{\Delta}^{\times}}g(\omega_{\Delta}^{i})\bigg{|}\;\geq\;1$ $\Rightarrow\prod_{i\in\mathbb{Z}_{\Delta}^{\times}}\bigg{|}\frac{g(\omega_{\Delta}^{i})}{\Delta^{k}}\bigg{|}\;\geq\;\left(\frac{1}{\Delta^{k}}\right)^{\varphi(\Delta)}$
$\Rightarrow\bigg{|}\prod_{i\in\mathbb{Z}_{\Delta}^{\times}}\frac{g(\omega_{\Delta}^{i})}{\Delta^{k}}\bigg{|}\;\geq\;\frac{1}{\Delta^{k\varphi(\Delta)}}.$
Now, the conjugate of $g(\omega_{\Delta}^{i})$, namely $\overline{g(\omega_{\Delta}^{i})}$, is nothing but $g(\omega_{\Delta}^{\Delta-i})$. Since $|g|_{1}\leq\Delta^{k}$, it follows that for any $i\in\mathbb{Z}_{\Delta}^{\times}$, $|g(\omega_{\Delta}^{i})/\Delta^{k}|\leq 1$. Therefore,
$\bigg{|}\prod_{i\in\mathbb{Z}_{\Delta}^{\times}}g(\omega_{\Delta}^{i})\bigg{|}\;\geq\;1$ $\Rightarrow\prod_{i\in\mathbb{Z}_{\Delta}^{\times}:i\leq\lfloor\frac{\Delta}{2}\rfloor}|g(\omega_{\Delta}^{i})|^{2}\;\geq\;1$
$\Rightarrow\prod_{i\in\mathbb{Z}_{\Delta}^{\times}:i\leq\lfloor\frac{\Delta}{2}\rfloor}\bigg{|}\frac{g(\omega_{\Delta}^{i})}{\Delta^{k}}\bigg{|}^{2}\;\geq\;(\frac{1}{\Delta^{k}})^{\varphi(\Delta)}$
$\Rightarrow\bigg{|}\frac{g(\omega_{\Delta})}{\Delta^{k}}\bigg{|}^{2}\;\geq\;\frac{1}{\Delta^{k\varphi(\Delta)}}$
$\Rightarrow\bigg{|}\frac{g(\omega_{\Delta})}{\Delta^{k}}\bigg{|}\;\geq\;\frac{1}{\Delta^{k\varphi(\Delta)/2}}.$ (3.9)
So, from Section 3.3, since $k=\lceil\log_{\Delta}2s_{f}\rceil$, we have
$\bigg{|}\frac{g(\omega_{\Delta})}{\Delta^{k}}\bigg{|}\;\geq\;\frac{1}{2^{\varphi(\Delta)/2}s_{f}^{\varphi(\Delta)/2}}.$
Remark 3.10.
The proof-technique of Theorem 1.5 gives the following. If the Fourier coefficients of a Boolean function $f$ are $k$ granular, then the Fourier coefficients of $f^{2}$ are $2k$-granular, and their absolute values are $\geq\frac{1}{2^{\varphi(\Delta)}s_{f}^{\varphi(\Delta)}}$.
3.4 Proof of Theorem 1.4
We have proved in Section 3.2 that if the Fourier weight of the function $f$ is concentrated on the $s$-largest Fourier coefficients, then $f$ is $\frac{\mu}{\sqrt{s}}$ close to a $k$-granular in the $\ell_{2}$ norm, where $k=\lceil\log_{\Delta}2s\rceil$. In this subsection, we will prove Theorem 1.4, which says that this function must be a Boolean valued function.
Proof.
Let us assume $k=\lceil\log_{\Delta}2s\rceil$. Let $B$ be the set of character functions corresponding to the $s$ largest coefficients and $S$ be set of the rest of the characters. Then, by Theorem 1.3, for all $\chi_{r_{i}}\in B$, $\widehat{f}(\chi_{r_{i}})$ is $\frac{\mu}{\sqrt{s}}$-close to $k$-granular. Therefore, for all $\chi_{r_{i}}\in B$, we can write $f(\chi_{r_{i}})$ as the sum of $\widehat{F}(\chi_{r_{i}})$ and $\widehat{G}(\chi_{r_{i}})$, that is,
@@ 3.3 Lower bound on the Fourier coefficients
where $|\widehat{F}(\chi_{r_{i}})|$ is $k$-granular and $|\widehat{G}(\chi_{r_{i}})|\leq\frac{\mu}{\sqrt{s}}$.
@@
Clearly, $F$ is $s$-sparse and $|\widehat{F}(\chi)|$ is $k$-granular for all Fourier coefficients $\widehat{F}(\chi)$ of $F$.
@@
$1=f^{2}=(F+G)^{2}=F^{2}+2FG+G^{2}=F^{2}+G(2F+G)=F^{2}+G(2f-G).$ (3.12)
Let H=G(2fG)H=G(2f-G). Let us look at the Fourier coefficients $\widehat{H}(\chi)$ of $H$.
[…]
@@
$\leq\frac{1}{2\times 2^{\varphi(\Delta)}s^{\varphi(\Delta)}}.$ (3.13)
Now, since $\widehat{F}(\chi)$ is $k$-granular, any Fourier coefficient $\widehat{F^{2}}(\chi)$ of $F^{2}$ is $2k$-granular. That is, $\widehat{F^{2}}(\chi)=\frac{g^{\prime}(\omega_{\Delta})}{\Delta^{2k}}$ and $g^{\prime}(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq\Delta-1$. Also, by Remark 3.10, $|\widehat{F^{2}}(\chi)|\geq\frac{1}{2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$.
Now, from Eq. 3.12, we have,
@@
where $\chi_{0}$ is the character which takes the value 11 at all points in $\mathcal{G}$, as defined in Lemma 2.4, part 1.
Since $\mid\widehat{F^{2}}(\chi)\mid\geq\frac{1}{2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$ and $\mid\widehat{H}(\chi)\mid<\frac{1}{2\times 2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$, these two cannot sum up to 0; therefore, $\widehat{F^{2}}(\chi)=0$, $\widehat{H}(\chi)=0$ for all $\chi\neq 0$.
@@
$\,\widehat{F^{2}}(\chi_{0})\,+\,\widehat{H}(\chi_{0})\;=\;1$
$\Rightarrow$ $\,\widehat{H}(\chi_{0})\;=\;1\,-\,\frac{g^{\prime}(\omega_{\Delta})}{\Delta^{2k}}$
$\Rightarrow$ $\,\widehat{H}(\chi_{0})\;=\;\frac{\Delta^{2k}\,-\,g^{\prime}(\omega_{\Delta})}{\Delta^{2k}}\;,$
which implies that $\widehat{H}(\chi_{0})$ is $2k$-granular if $\widehat{H}(\chi_{0})\neq 0$. So, by Remark 3.10, $|\widehat{H}(\chi_{0})|\geq\frac{1}{2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$, which is not possible since $|\widehat{H}(\chi_{0})|\leq\frac{1}{2\times 2^{\varphi(\Delta)}s^{\varphi(\Delta)}}$ by Section 3.4. So, $\widehat{H}(\chi_{0})=0$, hence $\widehat{F^{2}}(\chi_{0})=1$. Therefore $F^{2}(x)=0$ for all $x\in\mathcal{G}$, which implies that the range of $F$ is $\{-1,+1\}$. ∎
@@
4 A $\Delta$-independent lower bound
In this section, we discuss an interesting property of Boolean-valued functions from $\mathcal{G}$ to $\{-1,+1\}$, and eventually establish a $\Delta$-independent lower bound on the Fourier coefficient of an $s$-sparse Boolean-valued function.
For $x\in\mathcal{G}$ and $i\in\mathbb{Z}_{\Delta}^{\times}$, we define $i^{-1}x:=(i^{-1}x_{1},\ldots,i^{-1}x_{n})$, where $x=(x_{1},\ldots,x_{n})$ and $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{\Delta}^{\times}$ of $\mathbb{Z}_{\Delta}$.
Lemma 4.1.
Let $f:\mathcal{G}\to\{-1,+1\}$, and $\chi\in\operatorname{supp}(f)$. Then, for each $i\in\mathbb{Z}_{\Delta}^{\times}$, there exists a function $h_{i}:\mathcal{G}\to\{-1,+1\}$, defined by $h_{i}(x)=f(i^{-1}x)$ such that
@@ 4 A $\Delta$-independent lower bound
where $i^{-1}$ denotes the inverse of $i$ in the multiplicative group $\mathbb{Z}_{\Delta}^{\times}$ of $\mathbb{Z}_{\Delta}$.
Proof.
Proof of Part (1). Let $h_{i}(x):=f(i^{-1}x)$. Note that, by definition, $h_{i}$ is a Boolean-valued function. Further, by Definition 2.7:
$\hat{h_{i}}(\chi)$ $=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}h_{i}(x)\overline{\chi(x)}$
$=\frac{1}{|\mathcal{G}|}\sum_{ix\in\mathcal{G}}h_{i}(ix)\overline{\chi(ix)}$
$=\frac{1}{|\mathcal{G}|}\sum_{ix\in\mathcal{G}}f(x)\overline{\chi(x)^{i}}$
$=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}^{i}$
@@
Proof.
By Definition 2.7, we have $\hat{f}(\chi)=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}=\frac{u(\omega|\Delta)}{|\mathcal{G}|}$, where $u(\omega_{\Delta})=\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}$. Think of $u(X)\in\mathbb{Z}[X]$. Then, $\hat{h_{i}}(\chi)=\hat{f}(\chi^{i})=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}^{i}=\frac{u(\omega_{\Delta}^{i})}{|\mathcal{G}|}$; the last holds since
$u(\omega_{\Delta})\;=\;\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}\;\iff\;u(\omega_{\Delta}^{i})\;=\;\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}^{i}\;.$
Using Fact 2.35, we know that $u(\omega_{\Delta})\neq 0\iff u(\omega_{\Delta}^{i})\neq 0$. Hence, whenever $f(\chi)\neq 0,\ h_{i}(\chi)\neq 0$ as well, and the vice versa. Thus, the claim holds. ∎
The following corollary follows immediately from Lemma 4.1 (part 2).
Corollary 4.3.
@@
The following corollary follows immediately from Lemma 4.1.
Corollary $4.$
For all function $f:\mathcal{G}\to\{-1,+1\}$, $s_{f}\geq p-1$.
Proof.
From the proof of Lemma 4.1, it is clear that if $\chi\in\operatorname{supp}(f)$, then $\chi^{i}\in\operatorname{supp}(f)$ for all $i\in\mathbb{Z}_{\Delta}^{\times}$. Hence, sparsity of $f,\ s_{f}\geq|\mathbb{Z}_{\Delta}^{\times}|=\varphi(\Delta)$. ∎
Now, we prove a lower bound on the Fourier coefficient which is $\Delta$-independent; this is an immediate consequence of Corollary $4.$
Theorem 4.5 ($\Delta$-independent lower bound).
Let $f:\mathcal{G}\to\{-1,+1\}$ be a function. Let $\chi\in\operatorname{supp}(f)$. Then,
@@
Proof.
We know that $s_{f}\geq\varphi(\Delta)$, by Corollary $4.$ Hence, we combine this with Theorem 1.5 , to get
$|\widehat{f}(\chi)|\;\geq\;\frac{1}{2^{\varphi(\Delta)/2}s^{\varphi(\Delta)/2}}\;>\;\frac{1}{2^{s_{f}/2}s^{s_{f}/2}}.$
5 Sparse Boolean-valued function with small Fourier coefficients
In this section, for a fixed prime $p\geq 5$, and arbitrarily large $s$, we give an example of a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, such that the minimum of the absolute value of its Fourier coefficients is at most $o\left(1/s\right)$. More specifically, we give the details of the construction of Theorem 1.6.
Definition 5.1.
Lemma 5.2.
There is a Fourier coefficient of $\operatorname{\mathsf{AT}}$, whose absolute value is $\frac{1}{p^{nc}}$, where $c$ is a constant >1>1.
Proof.
@@ 5 Sparse Boolean-valued function with small Fourier coefficients
where we simply used the GP series with the ratio $\omega_{p}$. Note that
@@
$|\widehat{\operatorname{\mathsf{AT}}}(\chi_{2,2,\cdots 2})|\;=\;\bigg{(}\frac{1}{p\cos(\pi/p)}\bigg{)}^{n}\cdot\frac{1}{2^{n-1}}\;=\;2\bigg{(}\frac{1}{2p\cos(\pi/p)}\bigg{)}^{n}\;.$
@@
$\frac{1}{p^{nc}}\;=\;2\bigg{(}\frac{1}{2p\cos(\pi/p)}\bigg{)}^{n}\;\implies\;c=\log_{p}|2\cos(\pi/p)|+1-\frac{1}{n}\log_{p}2.$
Note that, for large enough $n$, and a fixed $p\geq 5$, $c>1$. Hence, the proof follows. ∎
Now we are ready to prove Theorem 1.6.
Proof of Theorem 1.6.
This directly follows, as we can claim from Lemma 5.2 that there exists a family of functions in $\mathbb{Z}_{p}^{n}$ whose absolute value of the minimum coefficient is not linear in $\frac{1}{\text{sparsity}}$, but actually $\Omega(\frac{1}{\text{sparsity}^{1+\epsilon_{p}}})$. where $\epsilon_{p}>0$, is a $p$-dependent constant. ∎
Lemma 5.3.
Proof.
@@
So comparing with Eq. 5.6, we can see that $c$ is bigger in this case. ∎
Remark 5.4.
When $p=3$, we know that $2\cos(\pi/3)=1$, and hence $c<1$ ($c\to 1$, when $n\to\infty$).
Remark $5.$
@@
$c=\log_{5}|(2\cos(\frac{\pi}{5}))|+1-\frac{1}{2}\cdot\log_{5}(2)=1.265$... (5.6)
@@
6 Testing $s$-sparsity
In this section our goal is to design an algorithm (see next page) that given a function $f:\mathcal{G}\to\{-1,+1\}$ accepts (with probability $\geq\frac{2}{3}$) if $f$ is $s$-sparse and rejects (with probability $\geq\frac{2}{3}$) if $f$ is $\epsilon$-far from any $s$-sparse Boolean function.
The algorithm 1’s primary concept closely aligns with that presented in [GOS+11]. Since we know from Theorem 1.5 that all the coefficients of an $s$-sparse function are more than $\frac{1}{2^{\varphi(\Delta)/2}s_{f}^{\varphi(\Delta)/2}}$, so the idea is to partition the Dual group (the set of characters) into $\Theta(s^{2})$ buckets, and for each bucket $B$ estimate the weight $\mathsf{wt}(B)$ of the bucket which is defined as follows:
Definition 6.1 (The weight function).
Let $f:\mathcal{G}\to\{-1,+1\}$ be a function. The weight of a bucket $B$ is defined by